a rectangle is four times longer than it is wide and the perimeter is 80m. what are the dimensions of the rectangle?
Answer by Alan3354(69443)

a rectangle is four times longer than it is wide and the perimeter is 80m. what are the dimensions of the rectangle?
===============
P = 2L + 2W = 80
L + W = 40
L = 5W
----
5W + W = 40
W = 20/3 meters
L = 100/3 meters